OpenStudy (anonymous):

To increase its labour force from 50 to 51 workers, a firm has to increase the daily wage rate from $600 to $610. What is the marginal cost of labour per day? A $10 B $510 C $610 D $1110

OpenStudy (anonymous):

as u knw marginal cost =change in total cost/change in labour

OpenStudy (anonymous):

Yup

OpenStudy (anonymous):

so ll be mrgc=($610-$600)/(51-50)

OpenStudy (anonymous):

So it would be 610-600/51-50 right? Which would be 10/1 right? Well, the answer isn't $10 :(

OpenStudy (anonymous):

That's what's bugging me.. the answer is actually $1110, and I don't know how to get to that.

OpenStudy (anonymous):

ll be $10/1 so mrgcst=$10

OpenStudy (anonymous):

It's an advanced level question... so it's not meant to be that easy, and the answer isn't that... so I don't know..

OpenStudy (hba):

\[Marginal \ cost =\frac{ Change \ i n \ total \ cost }{Change \ i n \ output }=\frac{ \Delta TC }{ \Delta q }\]

OpenStudy (anonymous):

Yes, thanks... but that doesn't help :(

OpenStudy (hba):

I know that :P

OpenStudy (hba):

Well i guess the labour is increasing by 1/> And it goes from 600-610 So, 10 $ Therfore, 10 /1=10

OpenStudy (anonymous):

The answer is $1110... I thought it was also $10...

OpenStudy (hba):

Are you sure it is $ 1110 ?

OpenStudy (anonymous):

Pretty sure, yeah, but not sure how to get to that answer...

OpenStudy (hba):

brb

OpenStudy (anonymous):

Sure

OpenStudy (anonymous):

its barely to knw

OpenStudy (hba):

Yuppie i got it :D

OpenStudy (anonymous):

How?!

OpenStudy (hba):

It is easy :P

OpenStudy (anonymous):

So say it!!!

OpenStudy (hba):

The wage rate is actually per labour So, Lets calculate 1=(600 * 50) Now lets calculate the new one. 2=(610*51) Now to calculate marginal cost Subtract them :) So, (610*51)-(600*50) =31,100-30,000=1,100 $

OpenStudy (anonymous):

Tricky examiners!!!! Come, on, you would have gone with $10 at first glance... that's how they want students to fail!!! Thanks for pointing it out though... and they weren't specific with it being one labour's wage xD How did you figure that out?

OpenStudy (hba):

My friend who is doing his BBA at LSE.

OpenStudy (anonymous):

What's BBA?

OpenStudy (hba):

Bachelor's in Business Administration.

OpenStudy (anonymous):

good ideal @hba

OpenStudy (hba):

Whaaaat :o

OpenStudy (anonymous):

thankx 4 assist

OpenStudy (hba):

Welcome :)

OpenStudy (anonymous):

Yeah, this helped a lot!!!. Still going to have to watch my back with examiners

OpenStudy (hba):

Best of luck with that :)

OpenStudy (phi):

I have not studied economics, but I would not overthink this problem 50 to 51 workers, a firm has to increase the daily wage rate from $600 to $610. What is the marginal cost of labour per day? the original cost per day is 50*600= 30000 the new cost per day is 51*610= 31110 So it looks like it costs $1110 per day to add 1 person

OpenStudy (hba):

^ Yeah that is correct :)